OpenStudy (anonymous):

What are the solutions to the system? (Hint: graphing is the easiest!) y = x2 + 2x – 45 y = -3x + 5

OpenStudy (anonymous):

would it be y=x(x+2)-45 -x^2-2x+y+45=0

OpenStudy (anonymous):

plug y=-3x+5 in the other eq. and find x

OpenStudy (anonymous):

Y=x2+2x-45 Y=0-3x+5 (multiply by -1) ------------- Y=x2+2x-45 -Y=0+3x-5 --------------- 0=x2+5x-50 Use baskara Delta=225 X1=5 X2=-10 Substitute these values on one equation Y=-3x+5 Y=-10 Y=35 So (5,-10) and (-10,35)

OpenStudy (anonymous):

\[x^2+2x-45=-3x+5\] \[x^2+2x+3x-45-5=0\] \[x^2+5x-45=0\] find x either by making factors or using quadratic formula

OpenStudy (anonymous):

correction write -50 in place of -45

OpenStudy (anonymous):

thank you both very much I don't know who I should medal but I will fan you both

OpenStudy (anonymous):

np